# FlowCast: Advancing Precipitation Nowcasting with Conditional Flow Matching

### v2

@@ 5 Conclusion
To ensure the reproducibility of our work, we have provided comprehensive supporting materials. Code: The full source code for our FlowCast model, including scripts for training and evaluation, is included in the supplementary material. The repository contains detailed instructions for setting up the required software environment and running the experiments. Datasets: Our work utilizes two datasets. The SEVIR dataset is a public benchmark, and details for access are provided by veillette2020sevir. The ARSO dataset was provided by the Slovenian Environment Agency (ARSO) for this research; we are actively collaborating with the agency to facilitate its public release in the near future. Experimental Details: Section 4.1 of the main paper provides a detailed description of our experimental setup. Further implementation details are available in Appendix A.1, including all model hyperparameters.
@@ A.1.1 VAE
Discriminator Architecture PatchGAN (isola2017image)
@@ A.1.2 FlowCast
The FlowCast architecture, adapted from Earthformer-UNet (gao2024prediff) for latent-space Conditional Flow Matching, has the following configuration:
@@
    • Base Feature Dimensionality: 196 units.
@@
        Generation: Generated with a channel multiplier of 4 relative to the base feature dimensionality (resulting in $196\times 4=784$ embedding channels).
@@ A.1.3 Evaluation Metrics
where $\Phi$ and $\phi$ denote the cumulative distribution function (CDF) and probability density function (PDF) of the standard normal distribution, respectively.

### v3

@@ Acknowledgments
To ensure the reproducibility of our work, we have provided comprehensive supporting materials. Code: The full source code for our FlowCast model, including scripts for training and evaluation, is available at https://github.com/b-rbmp/FlowCast. The repository contains detailed instructions for setting up the required software environment and running the experiments. Datasets: Our work utilizes two datasets. The SEVIR dataset is a public benchmark, and details for access are provided by Veillette et al. (2020). The ARSO dataset was provided by the Slovenian Environment Agency (ARSO) for this research; we are actively collaborating with the agency to facilitate its public release in the near future. Experimental Details: Section 4.1 of the main paper provides a detailed description of our experimental setup. Further implementation details are available in Appendix A.1, including all model hyperparameters.
@@ A.1.1 VAE
Discriminator Architecture PatchGAN (Isola et al., 2017)
@@ A.1.2 FlowCast
The FlowCast architecture, adapted from Earthformer-UNet (Gao et al., 2023) for latent-space Conditional Flow Matching, has the following configuration:
@@
    • Base Feature Dimensionality: 192 units.
@@
        Generation: Generated with a channel multiplier of 4 relative to the base feature dimensionality (resulting in $192\times 4=768$ embedding channels).
@@ A.1.3 Evaluation Metrics
where $\Phi$ and $\phi$ denote the cumulative distribution function (CDF) and probability density function (PDF) of the standard normal distribution, respectively. Following standard practice, we normalize the CRPS for each dataset by its maximum values, which scales our reported metrics by 1/255 for SEVIR and 1/57 for ARSO.